OpenStudy (anonymous):

Use basic identities to simplify to sines and cosines: (sinx/cot^2x)-(sinx/cos^2x)

OpenStudy (loser66):

your attempt?

OpenStudy (anonymous):

ive made several but i keep ending up with (sin^3-sin)/cos^2 but that doesnt look right...

OpenStudy (loser66):

\(tan =\dfrac{1}{cot}\rightarrow\)the first term is \(sin*tan^2\) ok?

OpenStudy (loser66):

tan =\(\dfrac{sin}{cos}\rightarrow\) the second term is \(\dfrac{1}{cos}*tan\)

OpenStudy (anonymous):

ohh i see i didnt split the second term into thatbut that makes sense

OpenStudy (anonymous):

i think i got it from ther thank you so much!

OpenStudy (loser66):

ok